OpenStudy (anonymous):

f(x)= -x^2 + 2x + 3 find the x intercepts

OpenStudy (solomonzelman):

set it equal to zero.\[-x^2 + 2x + 3=0\]everything times -1\[x^2 - 2x - 3=0\]\[(x+1)(x-3)=0\]\[x=-1~~~~~~~and~~~~~~~x=3\]

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

Anytime!